### v1

# The effects of dark energy on the matter-gravity coupling

Abstract
Dark energy perturbations are expected to modify the evolution of cosmological perturbations, producing observable effects on scalar perturbations and gravitational waves. We show that the effects of dark energy for a generic modified gravity model can be encoded in a momentum, space and polarization dependent effective matter-gravity coupling, in agreement with the results obtained in some specific theories, and that this is not just related to the effective Planck mass appearing in the action as a coefficient of the Ricci’s scalar. Using appropriately defined effective energy tensors, we show that the method can be extended to include also the effects of higher order perturbations. The generality of this effective description of the effects of dark energy on cosmological perturbations gives a solid theoretical justification for the phenomenological approaches often used in the analysis of cosmological observations, allowing to understand the mapping between the results of phenomenological analysis of observational data and theoretical models. The momentum and polarization dependence of the effective gravitational coupling implies that the gravitational energy emitted by a binary coalescence can have an additional frequency and polarization dependence, due to the effects of dark energy perturbations.
I Introduction
Dark energy accounts for most of the Universe energy density today, but its fundamental physical nature is still mysterious. For this reason modified gravity theories have been proposed as dark energy models, such as for example Horndeski Horndeski:1974wa. In these theories the field equations are modified with respect to general relativity (GR), and the new equations can be conveniently re-cast in the form of Einstein equations with a running Planck mass and effective stress-energy tensor associated to the new fields and geometrical terms, which is interpreted as dark energy.
This effective dark energy tensor has perturbations, which modify the GR cosmological perturbations equations, introducing new scale and time dependent effects. We show that these effects can be encoded in different effective time and momentum and polarization dependent matter-gravity couplings. As an application we discuss how this implies that the gravitational waves (GW) energy radiated by a binary system can have an additional frequency and polarization dependency, due to the interaction with dark energy.
II Modified gravity theories
In order to formulate a general theoretical framework let’s consider a theory defined by lagrangian of the form
$\mathcal{L}_{\rm{tot}}=\sqrt{-g}\,\left(\frac{1}{2}M_{eff}^{2}R+\rm L_{DE}+\rm L_{m}\right)\,$, (1)
where the subscripts $m$ and $\rm{DE}$ correspond to matter and dark energy, and $R$ is the Ricci’s scalar. The fields equations obtained by varying with respect to the metric are
$G_{\mu\nu}=\frac{1}{M_{\rm{eff}}^{2}}\left[T^{m}_{\mu\nu}+T^{DE}_{\mu\nu}\right]\,.$ (2)
where $T^{m}_{\mu\nu}$ and $T^{DE}_{\mu\nu}$ denote the matter and effective dark energy momentum tensors. The field equations can be always re-cast in this form by an appropriate manipulation. At this stage we have not specified the form of the dark energy Lagrangian, so eq.(1) represent a very large class of theories. In deriving the above equation we have assumed minimal coupling between the metric and the matter fields, i.e. we are using the Jordan frame.
III Cosmological effects on scalar perturbations
When studying the effects of dark energy on the background, the energy momentum tensors ar normally treated separately, but for cosmological perturbations the effects of dark energy are often modeled in terms of a modification of the general relativity (GR) equations, in terms of an effective gravitational coupling Ishak:2024jhs.
For example, in the Newton gauge the linearized scalar perturbations equation with respect to a FRW background takes the form
$\nabla^{2}\Psi=\frac{a^{2}}{2M^{2}_{\rm eff}}\left(\delta\rho_{m}+\delta\rho_{DE}\right)\,$,\\ (3)
[…]
For a given set of solutions of the cosmological perturbations equations we can re-write eq.(4) as
$\nabla^{2}\hat{\Psi}=\frac{a^{2}}{2M^{2}_{\rm eff}}\left(1+\frac{\delta\hat{\rho}_{DE}}{\delta\hat{\rho}_{m}}\right)\delta\hat{\rho}_{m}=\frac{1}{2}a^{2}G^{\Psi}_{\rm{eff}}\delta\hat{\rho}_{m}\\ \,$, (4)
where we are denoting with a hat the solutions of the perturbations equations, and we have defined the effective gravitational coupling as
$G^{\Psi}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta\hat{\rho}_{DE}}{\delta\hat{\rho}_{m}}\right)\,.$ (5)
See for example DeFelice:2011hq for an explicit calculation of $G^{\Psi}_{\rm{eff}}$, based on analytical solutions of the perturbations equations for Honrdeski theories. Note that since the cosmological perturbations solutions depends on both space and time, also $G^{\Psi}_{\rm{eff}}$ is both space and time dependent. Given the generality of the assumptions made to derive eq.(26), we can conclude that an effective Poisson equation of the form given in eq.(26), can provide a general effective description of the effects of dark energy on structure formation.
A similar procedure can be followed in Fourier space, giving an effective Poisson’s equation of the form
$\frac{k^{2}}{a^{2}}\Psi_{k}=\frac{1}{2}a^{2}G^{\Psi}_{\rm{eff}}(t,k)\delta\rho_{m,k}\\$ (6)
where we have now dropped the hat for ease of notation. This effective description allows to make the general prediction that the effective gravitational coupling can be both scale and time dependent, which can be useful for model independent analysis of observational data, treating $G^{\Psi}_{\rm{eff}}(t,k)$ as a phenomenological function constrained by data.
This effective approach requires to know the cosmological perturbations equations solutions to compute the effective gravitational coupling, i.e. it gives a general procedure to obtain the effective coupling, but it does not predict its explicit analytical form in terms of langrage coefficient. In this regard it is useful to use the effective field theory of dark energy to derive general analytical expression for $G^{\Psi}_{\rm{eff}}$, with the advantage of allowing to derive precise predictions for the interconnection with other observables Romano:2025apm; Romano:2025pcs.
A similar approach can also be applied to the other Bardeen potential $\Phi$, and for the equation relevant for weak lensing it gives
$\nabla^{2}(\Psi+\Phi)=a^{2}G^{\Psi+\Phi}_{\rm eff}\delta\rho_{m}\,\,$, (7)
which in momentum space takes the form
$\frac{k^{2}}{a^{2}}(\Psi_{k}+\Phi_{k})=a^{2}G^{\Psi+\Phi}_{\rm eff}(t,k)\delta\rho_{m,k}\,\,.$ (8)
Examples of explicit analytical calculations of $G^{\Psi+\Phi}$ and $G^{\Psi}$ for Horndeski theories in the deep sub-horizon limit, can be found in Linder:2015rcz. Note that this definition of effective gravitational coupling is consistent with the scale and time dependent ansatz used in the phenomenological analysis of large scale structure data Ishak:2024jhs.
IV Cosmological effects on gravitational waves
Dark energy also effects the propagation of gravitational waves, since the dark energy perturbations act as an additional effective energy-stress tensor. The perturbations of eq.(2) give rise to different modifications of the GR propagation equation. For general GWs $h_{ij}$ propagating in the $z$-direction Ezquiaga:2018btd, the spatial part of the metric corresponding to different polarizations can be written as
$h_{ij}=\begin{pmatrix}h_{S}+h_{+}&h_{\times}&h_{V1}\\ h_{\times}&h_{S}-h_{+}&h_{V2}\\ h_{V1}&h_{V2}&h_{L}\end{pmatrix}\,$, (9)
where $h_{+}$ and $h_{\times}$ are the tensor modes, $h_{V1,2}$ the vector polarizations, and $h_{S}$ and $h_{L}$ the transverse and longitudinal scalar modes. Ignoring the interaction of GWs with matter, the free action is given by Romano:2024apw; Gubitosi:2012hu
$\frac{a^{2}M^{2}_{\rm{eff}}}{c_{ij}^{2}}\Big[h^{\prime 2}_{A}-c_{ij}^{2}(\nabla h_{ij})^{2}\Big]$ (10)
giving the propagation equation
$h_{ij}^{\prime\prime}+2\mathcal{H}\Big(1-\frac{c_{ij}^{\prime}}{\mathcal{H}c_{ij}}+\frac{M_{\rm{eff}}^{\prime}}{\mathcal{H}M_{\rm{eff}}}\Big)h^{\prime}_{ij}-c_{ij}^{2}\nabla^{2}h_{ij}=0\,$, (11)
where $\mathcal{H}=a^{\prime}/a$. Including the interaction terms, we get
$h_{ij}^{\prime\prime}+2\mathcal{H}\Big(1-\frac{c_{ij}^{\prime}}{\mathcal{H}c_{ij}}+\frac{M_{\rm{eff}}^{\prime}}{\mathcal{H}M_{\rm{eff}}}\Big)h^{\prime}_{ij}-c_{ij}^{2}\nabla^{2}h_{ij}=\frac{a^{2}}{M^{2}_{\rm{eff}}}(\delta T^{m}_{ij}+\delta T^{DE}_{ij})=a^{2}G^{ij}_{\rm{eff}}\delta T^{m}_{ij}\,$, (12)
where we have defined the effective gravitational coupling for the different GWs polarizations as
$G^{ij}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta T^{DE}_{ij}}{\delta T^{m}_{ij}}\right)\,.$ (13)
Note that the indices $ij$ in $G^{ij}_{\rm{eff}}$ are not tensorial, but have only a labeling function, and are not summed in the source term or in any other equation. The full first order effective action corresponding to eq.(12), including the effects of interaction with matter, is
$\mathcal{L}^{(1)}_{h}=\frac{a^{2}M^{2}_{\rm{eff}}}{c_{ij}^{2}}\Big[h^{\prime 2}_{ij}-c_{ij}^{2}(\nabla h_{ij})^{2}\Big]+a^{4}G^{ij}_{\rm{eff}}h^{ij}\delta T^{m}_{ij}\,.$ (14)
Note that since cosmological perturbations are space and time dependent, in general also $G^{ij}_{\rm{eff}}(\eta,x)$ is expected to be space and time dependent. Contrary to GR, in a general modified gravity theory the vector and scalar degrees of freedom cannot be gauged away, because even in absence of matter perturbations the DE effective energy tensor will act as source for these polarizations. Note that the GWs propagation equation can be always re-cast in the form in eq.(12), by absorbing in $\delta T^{DE}_{ij}$ any modified gravity effect, including for example graviton mass terms. An alternative effective approach consists in encoding the effects of the source term in a frequency-time-polarization dependent effective speed Romano:2022jeh, which can be convenient for the calculation of the effects on gravity modification on the GW-EMW distance ratio.
In momentum space we can obtain a similar result
$h_{ij,k}^{\prime\prime}+2\mathcal{H}\Big(1-\frac{c_{ij}^{\prime}}{\mathcal{H}c_{ij}}+\frac{M_{\rm{eff}}^{\prime}}{\mathcal{H}M_{\rm{eff}}}\Big)h^{\prime}_{ij,k}+c_{ij}^{2}k^{2}h_{ij,k}=\frac{a^{2}}{M^{2}_{\rm{eff}}}(\delta T^{m}_{ij,k}+\delta T^{DE}_{ij,k})=a^{2}G^{ij}_{\rm{eff}}\delta T^{m}_{ij,k}\,$, (15)
where we have defined the scale dependent effective gravitational coupling for the different GWs polarizations as
$G^{ij}_{\rm{eff}}(\eta,k)=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta T^{DE}_{ij,k}}{\delta T^{m}_{ij,k}}\right)\,.$ (16)
V General relativity limit
When the dark energy perturbations are absent, i.e. when ${\delta T^{DE}_{ij}}=\delta\rho_{DE}=0$, the effective gravitational coupling for scalar perturbations and GWs are the same
$G^{\Psi+\phi}_{\rm{eff}}=G^{\Psi}_{\rm{eff}}=G^{ij}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\,$, (17)
and if the effective Planck mass is not running we obtain the GR limit
$G^{\Psi+\phi}_{\rm{eff}}=G^{\Psi}_{\rm{eff}}=G^{ij}_{\rm{eff}}=\frac{1}{M^{2}_{\rm p}}\,$, (18)
where $M_{p}$ is the Planck mass. Note that in general each GWs polarization and the different scalar perturbations have different effective gravitational couplings, and none of them is simply given by $1/{M^{2}_{\rm eff}}$.
VI Encoding higher order perturbations in the effective coupling
We have so far considered linear order metric perturbations, and at higher order $p$ other gauge invariant variables ${h}^{(p)}_{\mu\nu},\Psi^{p},\Phi^{p}$ can be defined Mollerach:2003nq; DeLuca:2019ufz; Chang:2020iji. Expanding the field equations (2) to higher orders in perturbations will give new equations which can always be re-cast in the canonical form Romano:2024apw
$\Box h^{(p)}_{ij}=\frac{a^{2}}{M^{2}_{\rm{eff}}}\left[\delta T^{m(p)}_{ij}+\delta T^{DE(p)}_{ij}\right]=a^{2}{G^{ij(p)}_{\rm{eff}}}\delta T^{m(p)}_{ij}\,$, (19)
where we have defined the effective gravitational coupling for the different GWs polarizations as
[…]
This can always be achieved by appropriately defining the effective perturbed energy-stress tensor $\delta T_{ij}^{DE(p)}$, even if the d’Alambert operator does not appear explicitly in the perturbed equations, by adding it on both sides of the equations. The d’Alambert operator is defined w.r.t. to the same background metric in all these equations, so they can be summed to give
$\Box h^{(N)}_{ij}=a^{2}{G^{ij(N)}_{\rm{eff}}}\delta T^{m(N)}_{ij}\,$, (21)
where we have used the fact that the d’Alambert operator is linear, and defined the summed GWs and energy-tensor perturbations
$h^{(N)}_{ij}=\sum^{N}_{p=1}h^{(p)}_{ij}\quad,\quad T^{(N)m}_{ij}=\sum^{N}_{p=1}\delta T^{(p)m}_{ij}\,$ (22)
and
$G^{ij(N)}_{\rm{eff}}=\left[\sum^{N}_{p=1}\frac{\delta T^{(p)m}_{ij}}{G^{ij(p)}_{\rm{eff}}}\right]^{-1}\delta T^{(N)m}_{ij}$ (23)
The quantities $h^{(N)}_{\mu\nu}$ are the physically observable GWs, obtained by summing the contributions from different perturbations orders. We have shown that the sum $h^{(N)}_{\mu\nu}$ of the solutions of the perturbations equations at different orders can be obtained as the solution of eq.(21) with an effective coupling given by eq.(23). This implies that eq.(21) can be used to model phenomenologically the effects of dark energy, including the non linear effects. The effective action corresponding to eq.(21) is
$\mathcal{L}^{(N)}_{h}=\frac{a^{2}M^{2}_{\rm{eff}}}{c_{ij}^{2}}\Big[{h_{ij}^{(N)^{\prime}}}^{2}-c_{ij}^{2}{\left(\nabla h_{ij}^{(N)}\right)}^{2}\Big]+a^{4}G^{ij(N)}_{\rm{eff}}h^{ij{(N)}}\delta T^{(N)m}_{ij}\,.$ (24)
For the modified Poisson’s equation we can obtain a similar result, by recasting the perturbed equations of order p in the canonical form
$\nabla^{2}\Psi^{(p)}=a^{2}G^{\Psi(p)}_{\rm{eff}}\delta{\rho}_{m}\,$,\\ (25)
where we have defined the effective gravitational coupling of order p as
$G^{\Psi(p)}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta\rho^{(p)}_{DE}}{\delta\rho^{(p)}_{m}}\right)\,$. (26)
Summing the equations at different orders in perturbations we get
$\nabla^{2}\Psi^{(N)}=a^{2}G^{\Psi(N)}_{\rm{eff}}\delta\rho^{(N)}_{m}\,$,\\ (27)
where we have defined the summed Bardeen potential $\Psi^{(N)}$ and energy-tensor perturbations
$\Psi^{(N)}=\sum^{N}_{p=1}\Psi^{(p)}\quad,\quad\delta\rho^{(N)}_{m}=\sum^{N}_{p=1}\delta\rho^{(p)}_{m}\,$, (28)
and the effective gravitational coupling is
$G^{\Psi(N)}_{\rm{eff}}=\left[\sum^{N}_{p=1}\frac{\delta\rho^{(p)}_{m}}{G^{\Psi(p)}_{\rm{eff}}}\right]^{-1}\delta\rho^{(N)}_{m}\,$. (29)
This shows that the sum $\Psi^{(N)}$ of the solutions of cosmological perturbations equations at different orders can be obtained as a solution of eq.(27), making it a convenient tool for the phenomenological study of the effects of dark energy.
VII Physical implications for scalar and tensor perturbations
We have shown that the effects of dark energy on cosmological perturbations can be modeled by a momentum, time and polarization dependent effective gravitational coupling. A direct implication of this general result is that the gravitational energy radiated by a binary system can acquire an additional dependence on the GWs frequency and polarization. This is due to the fact that the calculation of the radiated power involves the solution of the wave equation for a binary coalescence, and the momentum and polarization dependence of the effective gravitational coupling will induce new effects. Formally we can write the solution of the wave equation 19 at first order in perturbations as
$h_{ij}=\Box^{-1}\left({G^{ij}_{\rm{eff}}}\delta T^{m}_{ij}\right)$ (30)
which differ from the GR case in which $G^{ij}_{\rm{eff}}$ is a constant, and where we have assumed a Minkowski background. Since $G^{ij}_{\rm{eff}}$ is polarization and space dependent, different polarizations can have a different radiated power. Writing the wave equation in momentum space, the momentum dependence of $G^{ij}_{\rm{eff}}$ would also induce a new additional frequency dependency. In a general modified gravity theory there can be GWs of scalar and vector polarization, which cannot be gauged away, contrary to GR. Performing a scalar-vector-tensor decomposition of the metric perturbations, for the tensor part, i.e. the transverse and traceless (TT) piece Flanagan:2005yc, at leading order we obtain
$h^{TT}_{ij}=\frac{2G^{ij}_{\rm{eff}}}{r}\ddot{Q}^{TT}_{ij},$ (31)
where $Q$ is the traceless mass quadrupole moment. Accounting for the time and polarization dependence, but neglecting momentum dependence of the effective gravitational coupling, following a similar procedure as the one outlined in Wolf:2019hun we get the following expression for the radiated power of different polarizations
$P_{ij}=\frac{G^{ij}_{\rm{eff}}}{5}\left[\langle\dddot{Q}_{ij}\dddot{Q}^{ij}\rangle+\frac{2\dot{G}^{ij}_{\rm{eff}}}{G^{ij}_{\rm{eff}}}\langle\ddot{Q}_{ij}\dddot{Q}^{ij}\rangle+\left(\frac{\dot{G}^{ij}_{\rm{eff}}}{G^{ij}_{\rm{eff}}}\right)^{2}\langle\ddot{Q}_{ij}\ddot{Q}^{ij}\rangle\right].$ (32)
Since $G^{ij}_{\rm{eff}}$ depends on the GWs polarizations, the power radiated can also be polarization dependent, due to the potentially different way in which different polarizations interact with dark energy perturbations. We leave to a future work the derivation of a generalized quadrupole formula, including the frequency dependency due to the scale dependency of the effective gravitational coupling.
For scalar perturbations the scale dependence of the effective gravitational coupling allows to model phenomena such as the screening mechanism, i.e. a suppression the effects of dark energy on small scales, in order to pass local gravitational tests.
VIII Conclusions
We have shown that the effects of dark energy can be encoded in a momentum, space and polarization dependent effective gravitational coupling, and that this is not just related to the effective Planck mass appearing in the action as a coefficient of the Ricci’s scalar. The generality of this effective description of the effects of dark energy on cosmological perturbations gives a solid theoretical justification for the phenomenological approaches often used in the analysis of cosmological observations, and allows to understand the mapping between phenomenological analysis results Ishak:2024jhs; Nishizawa:2017nef and theoretical models. This effective description naturally predicts that the gravitational wave radiation emitted by binary systems can have an additional frequency and polarization dependence, and it will be interesting to test this prediction with future GWs observations.
For scalar perturbations the scale dependency of the effective gravitational coupling allows provide an effective description for the screening mechanism, i.e. the suppression of the effects of dark energy on small scales, which is often necessary in order to pass local gravitational tests.